Find all the multiples of 9 upto 100 ????????????????????

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

9 × 1 = 9
9 × 2 = 18
9 × 3 = 27
9 × 4 = 36
9 × 5 = 45
9 × 6 = 54
9 × 7 = 63
9 × 8 = 72
9 × 9 = 81
9 x 10 = 90
9 x 11 = 99
∴All the multiples of 9 upto 100 are 9, 18, 27, 36, 45, 54, 63, 72, 81, 90 and 99

Thanks